# Split-Post Microwave Displacement Transducer with Quadratic Readout

### v1

@@ I Introduction
Mechanical resonators come in diverse forms, including bulk acoustic wave (BAW) [14, 15, 16, 17, 18] and surface acoustic wave (SAW) devices [19], thin membranes [20, 21, 22, 23], and even lower-dimensional photonic structures [24, 25]. These can be integrated on-chip as micro-/nano-mechanical hybrid resonators coupled to microwave or optical photons [26, 27, 28]. The electromagnetic cavities themselves may be realised as lumped-element re-entrant modes [29, 30, 31, 32, 33, 34], whispering gallery modes [35, 36, 37, 38], or travelling-wave structures [10, 9, 11], including transmission line geometry [11]. Crucially, the opto-mechanical coupling between the microwave cavity mode and the mechanical displacement typically produces a linear coupling; however, higher-order dependencies of the microwave frequency on displacement can be engineered [39, 40, 41, 42, 43]. When cooled close to its motional ground state[44, 13], the mechanical mode enters the quantum regime [45], where its transitions obey Fermi’s golden rule [46]. Achieving ground-state cooling allows operation beyond the thermal noise floor [47, 22], and with sufficiently high cooperativity, a quadratic readout can effectively realise a multi-level quantum system [48, 39, 42, 40, 41]. Such systems exhibit phenomena including strong opto-mechanical coupling [32], parametric amplification [49, 23], ground-state cooling, and squeezing [28] and can store quantum information in entangled states [39, 48]. Macroscopic mechanical elements are powerful platforms for probing the classical–quantum boundary  [50, 32] and for sensing low-frequency signals, as their effective mass scales inversely with frequency. This makes them well suited to the kHz–MHz range, with applications to high-frequency gravitational-wave detection [51, 52, 53, 54, 55], resonant-mass gravitational antennas [56, 57], and dark-matter searches in the MHz–GHz band [58, 34], as well as precision tests of Lorentz symmetry [59, 50, 49] and the quantum nature of gravity [60, 46, 61, 62, 63, 64].
@@ V Conclusion
It was demonstrated that the split-post microwave cavity readout exhibited parametric quadratic behaviour due to the intrinsic symmetry of the microwave resonance. When the sapphire membrane was positioned at the centre or off-centre within the microwave resonator, the output response appeared quadratic or linear, respectively. This parametric readout system had the potential to be developed into a transducer capable of measuring energy when operated in its ground state. This work could be scaled to a single-graviton detection scheme, in which kilogram-scale detectors were impedance-matched to gram-scale detectors to measure energy deposited by a gravitational wave at kHz frequencies and thereby probe energy quantization. Such an approach, when enabled by ground-state cooling of the resonator, could yield promising results, particularly for phonon counting by building strongly coupled detectors capable of measuring energy [64, 46]. Our analysis further showed that the split-post resonator had a higher aspect ratio and greater field overlap in the region of the mechanical modes [32] than a single-post resonator [32, 31], resulting in a higher output coupling rate. By incorporating these features of split-post resonators, we were able to read out a non-piezoelectric membrane and characterise its response, contributing to the development of a quadratic transducer.

### v2

@@ I Introduction
Mechanical resonators come in diverse forms, including bulk acoustic wave (BAW) [14, 15, 16, 17, 18] and surface acoustic wave (SAW) devices [19], thin membranes [20, 21, 22, 23], and even lower-dimensional photonic structures [24, 25]. These can be integrated on-chip as micro-/nano-mechanical hybrid resonators coupled to microwave or optical photons [26, 27, 28]. The electromagnetic cavities themselves may be realised as lumped-element re-entrant modes [29, 30, 31, 32, 33, 34], whispering gallery modes [35, 36, 37, 38], or travelling-wave structures [10, 9, 11], including transmission line geometry [11]. Crucially, the opto-mechanical coupling between the microwave cavity mode and the mechanical displacement typically produces a linear coupling; however, higher-order dependencies of the microwave frequency on displacement can be engineered [39, 40, 41, 42, 43]. When cooled close to its motional ground state[44, 13], the mechanical mode enters the quantum regime [45], where its transitions obey Fermi’s golden rule [46, 47]. Achieving ground-state cooling allows operation beyond the thermal noise floor [48, 22], and with sufficiently high cooperativity, a quadratic readout can effectively realise a multi-level quantum system [49, 39, 42, 40, 41]. Such systems exhibit phenomena including strong opto-mechanical coupling [32], parametric amplification [50, 23], ground-state cooling, and squeezing [28] and can store quantum information in entangled states [39, 49]. Macroscopic mechanical elements are powerful platforms for probing the classical–quantum boundary  [51, 32] and for sensing low-frequency signals, as their effective mass scales inversely with frequency. This makes them well suited to the kHz–MHz range, with applications to high-frequency gravitational-wave detection [52, 53, 54, 55, 56], resonant-mass gravitational antennas [57, 58], and dark-matter searches in the MHz–GHz band [59, 34], as well as precision tests of Lorentz symmetry [60, 51, 50] and the quantum nature of gravity [61, 46, 62, 63, 64].
@@ V Conclusion
It was demonstrated that the split-post microwave cavity readout exhibited parametric quadratic behaviour due to the intrinsic symmetry of the microwave resonance. When the sapphire membrane was positioned at the centre or off-centre within the microwave resonator, the output response appeared quadratic or linear, respectively. This parametric readout system had the potential to be developed into a transducer capable of measuring energy when operated in its ground state. This work could be scaled to a single-graviton detection scheme, in which kilogram-scale detectors were impedance-matched to gram-scale detectors to measure energy deposited by a gravitational wave at kHz frequencies and thereby probe energy quantization. Such an approach, when enabled by ground-state cooling of the resonator, could yield promising results, particularly for phonon counting by building strongly coupled detectors capable of measuring energy [64, 46, 47]. Our analysis further showed that the split-post resonator had a higher aspect ratio and greater field overlap in the region of the mechanical modes [32] than a single-post resonator [32, 31], resulting in a higher output coupling rate. By incorporating these features of split-post resonators, we were able to read out a non-piezoelectric membrane and characterise its response, contributing to the development of a quadratic transducer.